Given the following construction project tasks and estimated durations: • Prepare foundation (3 days). • Assemble building (5 days). • Once the building is assembled, start the next two tasks at once: interior work (5 days) and landscaping (5 days). • Once the interior work is done, do the painting (2 days). • Once the landscaping is done, pave the driveway (1 day) • Once all tasks are completed, conduct the final inspection (1 day).

Prepare a PERT/CPM chart for this project and identify the overall project duration and critical path. Please charts are must a CPM and the PERT
Answer #1

Hey, here's the solution

Activity Symbol Activities Days
A Foundation              3
B Assembly Building              5
C Interior Work              5
D Landscaping              5
E Paiting              2
F Driveway              1
G Final Inspection              1

E-13 L=8 1-13 5DAYS 5DAYS E-15 E-16 L-16 YS L-15 1 DAY 8 3DAYS E=3 E=0 L=0 L=3 4 6 5DAYS 5 DAYS 1 DAY E-13 L-14 L=9 OVERALL P